The Ka for glycolic acid given by my textbook is 1.47x10^-4.Calculate the pH of a solution prepared by (This problem requires values in your textbooks specific appendices, which you ca
Answer #1

glycolic acid HO. VOH Chemical Formula: C2H,03 Molecular Weight: 76.05 (a) Number of moles of glycolic acid = mass/mol.wt. =

(b) Initial conc. M1 = 1.3412 M Initial volume = 35.0 mL Final volume = 350 mL final concentration = M2 = M1V1/V2 = 1.3412 *

(C) Initial conc. M1 = 0.13412 M Initial volume = 10.0 mL Final volume = 1.0 L = 1000 mL final concentration = M2 = M1V1/V2 =